Blossom Corp had sales of $300,000 in 2017. If management expects its sales to be $476,450 in 3 years, what is the rate at which the company’s sales are expected to grow? (If you solve this problem with algebra round intermediate calculations to 4 decimal places, in all cases round your final answer to 2 decimal places, e.g. 8.72%.)
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

476,450=300,000*(1+r/100)^3

(476,450/300,000)^(1/3)=(1+r/100)

(1+r/100)=1.1667

r=1.1667-1

=16.67%(Approx).